OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (lgbasallote):

perpendicular slope = -1/4 y+ 3 = -1/4(x-2) 4y + 12 = -x +2 x+ 4y +10 = 0